# Position: Spectral GNNs Are Neither Spectral Nor Superior for Node Classification

### 19 Mar 2026 v1

@@ Abstract
Spectral Graph Neural Networks (Spectral GNNs) for node classification promise frequency-domain filtering on graphs, yet rest on flawed foundations. Recent work shows that graph Laplacian eigenvectors do not in general have the key properties of a true Fourier basis, but leaves the empirical success of Spectral GNNs unexplained. We identify two theoretical glitches: (1) commonly used “graph Fourier bases” are not classical Fourier bases for graph signals; (2) (n1)(n-1)-degree polynomial polynomials ($n$ = number of nodes) can exactly interpolate any spectral response via a Vandermonde system, so the usual “polynomial approximation” narrative is not theoretically justified. The effectiveness of GCN is commonly attributed to spectral low-pass filtering, yet we prove that low- and high-pass behaviors arise solely from message-passing dynamics rather than Graph Fourier Transform–based spectral formulations. We then analyze two representative directed spectral models, MagNet and HoloNet. Their reported effectiveness is not spectral: it arises from implementation issues that reduce them to powerful MPNNs. When implemented consistently with the claimed spectral algorithms, performance becomes weak. This position paper argues that: for node classification, Spectral GNNs neither meaningfully capture the graph spectrum nor reliably improve performance; competitive results are better explained by their equivalence to MPNNs, sometimes aided by implementations inconsistent with their intended design.
@@ 3.2 Pitfalls of Truncated Polynomial Approximations
Once $\bm{\alpha}$ is fixed, each coefficient $\alpha_{k}$-whether multiplying $\lambda^{0}$ or $\lambda^{n-1}$-contributes equally to matching $\bm{\theta}^{*}$ on the spectrum. This reveals that the recent surge of truncated polynomial bases inadvertently favors low-degree coefficients $\alpha_{0}$, $\alpha_{1}$, … (corresponding to local neighborhood aggregation) under-utilizing higher-degree coefficients.
Simply proposing new polynomial families—such as Chebyshev expansions (Defferrard et al., 2016), first-order Chebyshev expansion (Kipf and Welling, 2016), Bernstein bases (He et al., 2021), or others (Xu et al., 2024)—amounts to blindly exploring different values of $\alpha_{k}$. For example, ChebNet (Defferrard et al., 2016), effectively sets $\alpha_{0}$=1, $\alpha_{1}$=1, while GCN (Kipf and Welling, 2016) corresponds to $\alpha_{0}$=2, $\alpha_{1}$=-1 (after renormalization). These remain particular solutions within the full interpolation space, not structurally privileged choices.
Furthermore, Equation 3 is inherently dataset-specific: $\bm{\Lambda}$ and $\bm{\theta}$ vary across graphs. Spectral GNNs thus imply dataset-specific models, undermining their claimed generality.
@@ 3.3 What Polynomial Approximation Truly Leads To
The two conceptual flaws identified above — the graph Fourier basis interpretation (Guo et al., 2025) and the polynomial approximation pitfalls — turn out, in combination, to be accidentally productive. With their combination, spectral filters are polynomial functions of aggregation via graph Laplacian; and since the Laplacian encodes local adjacency, different polynomial filters correspond to different weighted combinations of spatial aggregations, with polynomial order directly controlling the maximum hop distance (Balcilar et al., 2021). Two conceptually flawed motivations thus converge on a correct and interpretable computational primitive: the MPNN.
@@ 6 Conclusion
In summary, spectral GNNs provide no unique theoretical or practical benefits. We advocate that future research in graph learning should be grounded in spatial-domain frameworks that are conceptually transparent and theoretically robust, rather than in spectral assumptions that lack rigorous justification.
@@ 7 Alternative Views
First, as discussed in Sections 2.1 and 3.1, spectral GNNs are typically built from two key ingredients: (i) graph Fourier basis and (ii) polynomial approximations. For directed graphs, a common strategy is to construct a Hermitian operator (Section 5.1.1) so that a Fourier-type basis remains available.
@@
Third, spectral GNNs have been extended to directed graphs, and several widely used directed spectral architectures—such as MagNet (Zhang et al., 2021) and HoloNet (Koke and Cremers, 2024)—have been highly influential, directly and indirectly shaping many follow-up works. More broadly, new variants of spectral GNNs continue to appear at scale each year, reflecting sustained interest and optimism within the community.

### 26 Mar 2026 v2

@@ Abstract
Spectral Graph Neural Networks (Spectral GNNs) for node classification promise frequency-domain filtering on graphs, yet rest on flawed foundations. Recent work shows that graph Laplacian eigenvectors do not in general have the key properties of a true Fourier basis, but leaves the empirical success of Spectral GNNs unexplained. We identify two theoretical glitches: (1) commonly used “graph Fourier bases” are not classical Fourier bases for graph signals; (2) (n1)(n-1)-degree polynomials ($n$ = number of nodes) can exactly interpolate any spectral response via a Vandermonde system, so the usual “polynomial approximation” narrative is not theoretically justified. The effectiveness of GCN is commonly attributed to spectral low-pass filtering, yet we prove that low- and high-pass behaviors arise solely from message-passing dynamics rather than Graph Fourier Transform–based spectral formulations. We then analyze two representative directed spectral models, MagNet and HoloNet. Their reported effectiveness is not spectral: it arises from implementation issues that reduce them to powerful MPNNs. When implemented consistently with the claimed spectral algorithms, performance becomes weak. This position paper argues that: for node classification, Spectral GNNs neither meaningfully capture the graph spectrum nor reliably improve performance; competitive results are better explained by their equivalence to MPNNs, sometimes aided by implementations inconsistent with their intended design.
@@ 3.2 Pitfalls of Truncated Polynomial Approximations
Crucially, this constitutes a finite interpolation problem rather than a traditional polynomial approximation task. Classical approximation techniques—such as Taylor or Chebyshev expansions—progressively improve as terms of increasing degree refine the estimate of a continuous function $f$. In contrast, interpolation over a Vandermonde system lacks this hierarchical structure: all polynomial orders are equally important. Once the coefficients $\bm{\alpha}$ are determined, each coefficient $\alpha_{k}$-regardless of degree, whether multiplying $\lambda^{0}$ or $\lambda^{n-1}$-contributes equally to matching $\bm{\theta}^{*}$ on the spectrum.
However, empirical results in previous studies, such as Jiang et al. (2025), are inconsistent with this theoretical deduction. Experiments consistently show that aggregating from low-order neighborhoods outperforms aggregation from higher-order ones — the opposite of what the spectral motivation would suggest. This inconsistency calls into question the validity of the theoretical framework from which the deduction was derived, and provides further evidence that the Graph Fourier Transform is a flawed foundation for modeling graph signals. This theoretical deduction also sits in tension with a recurring empirical trend: truncated polynomial approximations inadvertently favour low-degree coefficients $\alpha_{0}$, $\alpha_{1}$, … (corresponding to local neighborhood aggregation), while under-utilizing higher-degree terms.
We now follow further deductions stemming from this flawed premise. Spectral GNNs that propose alternative polynomial families, such as Chebyshev expansions (Defferrard et al., 2016), first-order Chebyshev expansion (Kipf and Welling, 2016), Bernstein bases (He et al., 2021), or others (Xu et al., 2024), amounts to blindly exploring different values of $\alpha_{k}$. For example, ChebNet (Defferrard et al., 2016), effectively sets $\alpha_{0}$=1, $\alpha_{1}$=1, while GCN (Kipf and Welling, 2016) corresponds to $\alpha_{0}$=2, $\alpha_{1}$=-1 (after renormalization). These are merely particular points in the overall interpolation space, without any structurally privileged status. Furthermore, Equation 3 is inherently dataset-specific: $\bm{\Lambda}$ and $\bm{\theta}$ vary across graphs. Spectral GNNs thus imply dataset-specific models, undermining their claimed generality. Collectively, these deductions highlight an internal inconsistency: conclusions derived from the theoretical formulation of spectral GNNs contradict their empirical practice.
In summary, we expose the limitations of spectral GNNs from multiple perspectives: (1) each of the two constituent steps is independently problematic, rooted in a misapplication of Fourier analysis and polynomial approximation respectively. (2) theoretical implications contradict empirical evidence that low-order terms are more important than higher-order terms; and (3) the design choices of spectral GNNs, such as favouring low-order terms, unprincipled parameter search, and implicit dataset-specificity that undermines their claimed generality, are inconsistent with their own theoretical deductions.
@@ 3.3 What Polynomial Approximation Truly Leads To
The two conceptual flaws identified above — the graph Fourier basis interpretation (Guo et al., 2025) and the polynomial approximation pitfalls — turn out, in combination, to be accidentally productive. With their combination, spectral filters are polynomial functions of aggregation via graph Laplacian; and since the Laplacian encodes local adjacency, different polynomial filters correspond to different weighted combinations of spatial aggregations, with polynomial order directly controlling the maximum hop distance (Balcilar et al., 2021). Thus, although each ingredient is internally flawed, together they converge on a correct and interpretable computational primitive: the MPNN.
@@ 6 Conclusion
In summary, Spectral GNNs provide no unique theoretical or practical benefits. Spectral GNNs rest on two theoretically flawed steps that, in combination, accidentally recover the MPNN framework — lending them an effectiveness their spectral justification does not warrant.
@@ 7 Alternative Views
First, as discussed in Sections 2.1 and 3.1, Spectral GNNs are typically built from two key ingredients: (i) graph Fourier basis and (ii) polynomial approximations. For directed graphs, a common strategy is to construct a Hermitian operator (Section 5.1.1) so that a graph Fourier-type basis remains available.
@@
Third, Spectral GNNs have been extended to directed graphs, and several widely used directed spectral architectures, such as MagNet (Zhang et al., 2021) and HoloNet (Koke and Cremers, 2024), have been highly influential, directly and indirectly shaping many follow-up works. More broadly, new variants of Spectral GNNs continue to appear at scale each year, reflecting sustained interest and optimism within the community.